If 5.0 g of KClO3 are decomposed, what volume of O2 is produced at STP?


1
Expert's answer
2021-03-04T06:35:54-0500

Solution:

The balanced chemical equation:

2KClO3 → 2KCl + 3O2

According to the equation above: n(KClO3) / 2 = n(O2) / 3


Moles of KClO3 = Mass of KClO3 / Molar mass of KClO3

The molar mass of KClO3 is 122.55 g mol-1.

Hence,

n(KClO3) = (5.0 g) / (122.55 g mol-1) = 0.0408 mol


n(O2) = 3 × n(KClO3) / 2 = (3 × 0.0408 mol) / 2 = 0.0612 mol (according to the equation)


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.0612 mol of O2 occupies:

(0.0612 mol × 22.4 L) / 1 mol = 1.37088 L of O2 = 1.37 L of O2

At STP, 0.0612 mol of O2 occupies a volume equal to 1.37 L.


OR (according to the equation):

(5.0 g KClO3)×(1 mol KClO3 / 122.55 g KClO3)×(3 mol O2 / 2 mol KClO3)×(22.4 L O2 / 1 mol O2) = 1.37 L

1.37 L O2


Answer: 1.37 liters of O2 is produced at STP.